# 2INFN, Laboratori Nazionali del Gran Sasso, I-67100 Assergi, Italy 3Universität Heidelberg, Zentrum für Astronomie (ZAH), Institut für Theoretische Astrophysik, Albert Ueberle Str. 2, 69120, Heidelberg, Germany 4Universität Heidelberg, Interdisziplinäres Zentrum für Wissenschaftliches Rechnen, Heidelberg, Germany

### v1

@@ A.2 Mass ratio and secondary mass
    The $q$ distribution from Sana et al. (2012), $\xi(q)\propto q^{-0.1}$, with $q\in[0.1,1]$, which fits the mass ratios of O- and B-type binaries in the local Universe; here we set $m_{\mathrm{ZAMS,2}}\geq 2.2\,{\rm M}_{\odot}$.
@@
    The $q$ distribution from Stacy & Bromm (2013), $\xi(q)\propto q^{-0.55}$, with $q\in[0.1,1]$. This distribution was obtained by fitting Pop. III stars generated in cosmological simulations. Here we assume $m_{\mathrm{ZAMS,2}}\geq 2.2\,{\rm M}_{\odot}$.
@@ A.3 Orbital period
    $\xi(\Pi)\propto\Pi^{-0.55}$, with $\Pi=\log{(P/\mathrm{day})}\in[0.15,5.5]$ from Sana et al. (2012).
@@
    $\xi(\Pi)\propto\exp{\left[-\left(\Pi-\mu\right)^{2}/\left(2\sigma^{2}\right)\right]}$, that is a Gaussian distribution with $\mu=5.5$ and $\sigma=0.85$ (Stacy & Bromm 2013).
@@ A.4 Eccentricity
    $\xi(e)\propto e^{-0.42}$ with $e\in[0,1)$ (Sana et al. 2012).
@@
    A thermal distribution $\xi(e)\propto e$ with $e\in[0,1)$ (Heggie 1975; Tanikawa et al. 2021b; Hartwig et al. 2016; Kinugawa et al. 2014), which favors highly eccentric systems. As shown by Park et al. (2021, 2023), Pop. III binaries form preferentially with high orbital eccentricity.
@@
In Figs. 3 and 8, we compare the MRDs of BHNSs and BNSs from Pop. III and Pop. II stars with those from Pop. II-I. These MRDs were derived by Iorio et al. (2023) using a fiducial configuration. Below, we summarize the main properties of this model and describe the computation of the MRD. For further details, we refer the reader to Iorio et al. (2023).
@@
The resulting BHNS and BNS merger catalogs are used as input for $\mathcal{R}$ate. To compute the MRD of these systems, we use Eq. 2 from Sec. 2.7, assuming the star formation rate density $\psi(z)$ from Madau & Fragos (2017):
@@
where we assume the metallicity distribution $p(z^{\prime},Z)$ from Madau & Fragos (2017):
@@
Here, $\langle\log Z(z^{\prime})/\mathrm{Z_{\odot}}\rangle=\log\langle Z(z^{\prime})/\mathrm{Z_{\odot}}\rangle-\ln(10)\,\sigma_{\rm Z}^{2}/2$, and we adopt $\sigma_{\rm Z}=0.2$ (Bouffanais et al. 2021).
@@
Appendix C Model selection
Here, we briefly summarize the method presented by Mould et al. (2023). According to Bayes’ theorem, the posterior distribution of the measured properties of a GW event is given by
@@ Appendix C Model selection
where $\mathcal{L}(\@vec{d}|\@vec{\theta})$ is the likelihood that the vector of waveform parameters ${\theta}$ (e.g., masses, spins, redshifts, etc.) produces the observed data ${d}$, and $\pi(\@vec{\theta}|U)$ is the uninformative prior on the parameters. The normalization factor,
[…]
@@
Integrating over ${\theta}$ and rearranging, we obtain the Bayes factor, which quantifies the relative likelihood of the two prior models:
[…]
A value of $\mathcal{B}_{A/U}>1$ ($<1$) implies that the astrophysical model $A$ is more (less) likely than the uninformative prior $U$ to have produced the observed data ${d}$. A decisive conclusion is typically reached when $|\ln\mathcal{B}_{A/U}|\gtrsim 5$ (Jeffreys 1939). If we consider two different informative models, $A$ and $B$, the Bayes factor is simply $\mathcal{B}_{A/B}=\mathcal{B}_{A/U}/\mathcal{B}_{B/U}$.

### v2

@@ A.2 Mass ratio and secondary mass
    The $q$ distribution from sana2012, $\xi(q)\propto q^{-0.1}$, with $q\in[0.1,1]$, which fits the mass ratios of O- and B-type binaries in the local Universe; here we set $m_{\mathrm{ZAMS,2}}\geq 2.2\,{\rm M}_{\odot}$.
@@
    The $q$ distribution from sb2013, $\xi(q)\propto q^{-0.55}$, with $q\in[0.1,1]$. This distribution was obtained by fitting Pop. III stars generated in cosmological simulations. Here we assume $m_{\mathrm{ZAMS,2}}\geq 2.2\,{\rm M}_{\odot}$.
@@ A.3 Orbital period
    $\xi(\Pi)\propto\Pi^{-0.55}$, with $\Pi=\log{(P/\mathrm{day})}\in[0.15,5.5]$ from sana2012.
@@
    $\xi(\Pi)\propto\exp{\left[-\left(\Pi-\mu\right)^{2}/\left(2\sigma^{2}\right)\right]}$, that is a Gaussian distribution with $\mu=5.5$ and $\sigma=0.85$ (sb2013).
@@ A.4 Eccentricity
    $\xi(e)\propto e^{-0.42}$ with $e\in[0,1)$ (sana2012).
@@
    A thermal distribution $\xi(e)\propto e$ with $e\in[0,1)$ (heggie1975; tanikawa2021; hartwig2016; kinugawa2014), which favors highly eccentric systems. As has been shown by park2021; park2023, Pop. III binaries form preferentially with high orbital eccentricity.
@@
In Figs. 3 and 8, we compare the MRDs of BHNSs and BNSs from Pop. III and Pop. II stars with those from Pop. II-I. These MRDs were derived by sevn2023 using a fiducial configuration. Below, we summarize the main properties of this model and describe the computation of the MRD. For further details, we refer the reader to sevn2023.
One million binary star systems were simulated at 15 different metallicities, $Z\in[10^{-4},\,3\times 10^{-2}]$. The initial conditions for these simulations match those of the kro1 model, except for the mass range considered. In sevn2023, the primary mass $m_{\rm ZAMS,1}$ is drawn from a Kroupa IMF (kroupa2001) in the range $5,150]\,{\rm M}_{\odot}$, as opposed to $2,600]\,{\rm M}_{\odot}$ in kro1. The initial distributions of mass ratio and orbital period follow sana2012. The minimum secondary mass is set to $m_{\rm ZAMS,2}=2.2\,{\rm M}_{\odot}$. The treatment of mass transfer, compact object formation, SNe, pair-instability SNe, and natal kicks is the same as described in Sec. 2.
@@
where we assume the metallicity distribution $p(z^{\prime},Z)$ from madau2017:
@@
Here, $\langle\log Z(z^{\prime})/\mathrm{Z_{\odot}}\rangle=\log\langle Z(z^{\prime})/\mathrm{Z_{\odot}}\rangle-\ln(10)\,\sigma_{\rm Z}^{2}/2$, and we adopt $\sigma_{\rm Z}=0.2$ (bouffanais2021).
@@
Appendix C Impact of Pop. III star-formation rate models on BHNS and BNS MRDs
In this Section, we evaluate how the choice of the Pop. III star formation rate density model affects the results presented in Sec. 3.1.1 and Sec. 3.2.1. Specifically, we focus on the Pop. III models that yield the highest MRDs in Fig. 3 and Fig. 8 - lar1 and kro1, respectively - and recompute the MRDs assuming star formation rate densities from hartwig2022, liu2020, skinner2020, jaacks2019, and desouza2011. It should be noted that the optimistic star-formation rate density model presented by desouza2011 is in tension with the cosmological parameters from aghanim2020. As a consequence, dS11 is a rescaled version of it by 0.3, following the formalism presented in kinugawa2019.
Figure 12 shows the MRDs for BHNS systems. The H22, LB20, and J19 models yield MRDs that remain below the MRD from Pop. II BHNSs up to high redshifts ($z\in\left[17,\,23\right]$). The SW20 model yields the lowest BHNS MRD of all the considered star-formation rate models for Pop. III stars. In contrast, the dS11 model produces a BHNS MRD slightly higher than the Pop. II one at $z\lesssim 7$ and then again from $z\sim 22$ onward. Across the different models for the Pop. III star formation rate evolution, the MRDs peak at $\mathcal{R}_{\rm BHNS}\in\left[2,\,20\right]\,\rm{Gpc^{-3}\,yr^{-1}}$. At $z=0$, the MRDs are consistently comparable to, or greater than, the fiducial value ($\sim 0.03\,\rm Gpc^{-3}\,yr^{-1}$), implying that the MRD of Pop. III BHNSs with $m_{1}>20\,{\rm M}_{\odot}$ is always equal to or higher than that of BHNSs from Pop. II stars. This outcome highlights the robustness of our findings from the fiducial model.
Figure 13 shows the MRDs of BNSs for the various Pop. III star formation rate density models. The H22, SW20, and J19 models exceed the Pop. II MRD only from intermediate redshifts ($z\in\left[7,\,11\right]$) while LB20 and dS11 remain dominant throughout the entire redshift range. At $z=0$, the LB20 model yields the highest MRD. As in Fig. 8, a sharp decline is visible at $z<5$, driven by the combination of short delay times and star formation rate densities that do not extend to lower redshifts; the sole exception is LB20, because liu2020 extrapolated their Pop. III star formation rate density down to $z=0$.
Figure 12: BHNS MRDs for the model producing the largest MRD in Fig. 3 (lar1), considering different prescription for the Pop. III star formation rate density. Thick blue line: hartwig2022; violet dotted line: liu2020; magenta dashed line: skinner2020; orange dash-dotted line: jaacks2019; yellow dash-dot-dotted line: rescaled desouza2011. As a comparison we also plot the BHNS MRD for early Pop. II (gray line) and Pop. II-I BHNSs (black line).
Figure 13: Same as Fig. 12 but for the BNS MRDs, considering the Pop. III model producing the largest MRD in Fig. 8 (kro1).
Appendix D Model selection
Here, we briefly summarize the method presented by mould2023. According to Bayes’ theorem, the posterior distribution of the measured properties of a GW event is given by
@@ Appendix D Model selection
where $\mathcal{L}(\@vec{d}|\@vec{\theta})$ is the likelihood that the vector of waveform parameters $\@vec{\theta}$ (e.g., masses, spins, redshifts, etc.) produces the observed data d\@vec{d}, and $\pi(\@vec{\theta}|U)$ is the uninformative prior on the parameters. The normalization factor,
$\mathcal{Z}(\@vec{d}|U)=\int\mathcal{L}(\@vec{d}|\@vec{\theta^{\prime}})\,\pi(\@vec{\theta^{\prime}}|U)\,d\@vec{\theta^{\prime}}$ (8)
@@
Integrating over $\@vec{\theta}$ and rearranging, we obtain the Bayes factor, which quantifies the relative likelihood of the two prior models:
$\mathcal{B}_{A/U}=\frac{\mathcal{Z}(\@vec{d}|A)}{\mathcal{Z}(\@vec{d}|U)}=\int p(\@vec{\theta}|\@vec{d},U)\,\frac{\pi(\@vec{\theta}|A)}{\pi(\@vec{\theta}|U)}\,d\@vec{\theta}.$ (10)
A value of $\mathcal{B}_{A/U}>1$ (<1<1) implies that the astrophysical model $A$ is more (less) likely than the uninformative prior $U$ to have produced the observed data d\@vec{d}. A decisive conclusion is typically reached when $|\ln\mathcal{B}_{A/U}|\gtrsim 5$ (jeffreys1939). If we consider two different informative models, $A$ and $B$, the Bayes factor is simply $\mathcal{B}_{A/B}=\mathcal{B}_{A/U}/\mathcal{B}_{B/U}$.